the sum of magnitudes of two forces acting at a point is 18N and the magnitude of their resultant is 12N. if the resultant makes an angle of 90 degree with the force of smaller magnitude find the magnitudes of the two forces.

Grade:11

1 Answers

Vikas TU
7 years ago
let a and b ve the forces whose sum is given as:
a + b = 18.....................................(1)
root(a^2 + b^2 + 2abcosthetha) = 12.........................(2)
cos90 = R.b/|R||b|
.i.e.
R.b = 0 
(a+b).b = 0 
or
a.b = -|b|^2................(3)
use these 3 eqns. and solve for a,b, and costhetha.